# Observability rank conditions for analysing practical identifiability a priori

### v1

@@ 3.4 Output derivatives
To support this claim, we perform the following numerical experiment. We generate synthetic data by simulating model $\mathcal{M}_{3.4}$ from five different experimental conditions, adding 10% of Gaussian noise. This level of noise is one that we could reasonably expect in many practical situations. Figure 2 shows in its left side the noiseless signal along with the corresponding noisy pseudo-experimental data of one of the experiments; the other four are similar. Then, we try to estimate the values of the parameters and of the initial conditions of each experiment. The right hand side of 2 shows the profile likelihood obtained for $\theta_{1}$. The parameter is clearly identifiable, and the plot exhibits an optimum at the nominal value ($\theta_{1}=1$).
Figure 2: Model $\mathcal{M}_{3.4}$. The panel on the left shows the noisy data (blue circles) of one of the five experiments used for parameter estimation, as well as the noiseless simulation of the output (red line). The panel on the right shows the resulting profile likelihood of parameter $\theta_{1}$. It can be seen that the parameter is practically identifiable.
3.5 Relative identifiability and observability
@@
The idea is to perform a singular value decomposition (SVD) of the numerical $\mathcal{O}_{I}$. If the model is practically observable, all the singular values will be of similar magnitude, i.e. the condition number (the ratio of the largest singular value over the smallest one) will be small Dochain et al. (1997). Likewise, the smallest singular value indicates the worst estimate, so ideally it should large. The idea of analysing a linear system in this way was brought to the attention of the systems and control community by Moore (1981).

### v2

@@ 3.4 Output derivatives
To support this claim, we perform the following numerical experiment. We generate synthetic data by simulating model $\mathcal{M}_{3.4}$ from five different experimental conditions, adding Gaussian noise with zero mean and a standard deviation of 2.0. This level of noise is one that we could reasonably expect in many practical situations. Figure 2 shows in its left side the noiseless signal along with the corresponding noisy pseudo-experimental data of one of the experiments; the other four are similar. Then, we try to estimate the values of the parameters and of the initial conditions of each experiment. The right hand side of Figure 2 shows the bootstrap histogram obtained for $\theta_{1}$, which shows that $\theta_{1}$ is clearly identifiable, with median and mean estimates almost identical to the nominal value ($\theta_{1}=1$) and a very low standard deviation.
Figure 2: Model $\mathcal{M}_{3.4}$. The panel on the left shows the noisy data (blue circles) of one of the five experiments used for parameter estimation, as well as the noiseless simulation of the output (red line). The panel on the right shows the resulting bootstrap of parameter $\theta_{1}$. It can be seen that the parameter is practically identifiable.
3.5 Numerical identifiability and observability through singular value decomposition
@@
The idea is to perform a singular value decomposition (SVD) of the numerical $\mathcal{O}_{I}$. If the model is practically observable, all the singular values will be of similar magnitude, i.e. the condition number (the ratio of the largest singular value over the smallest one) will be small Dochain et al. (1997). Likewise, the smallest singular value indicates the worst estimate, so ideally it should be large. The idea of analysing a linear system in this way was brought to the attention of the systems and control community by Moore (1981). Stigter and Joubert (2021) used a similar approach, but performing SVD of a sensitivity matrix instead of an observability-identifiability one.
To illustrate this methodology, we will recall the previously considered examples $\mathcal{M}_{3.2}$ and $\mathcal{M}_{3.4}$, as well as the following model:
$\mathcal{M}_{3.5}:\left\{\begin{array}[]{ll}\dot{G}&=u-(c+s_{i}\cdot I)\cdot G,\\ \dot{\beta}&=\beta\cdot\left(\dfrac{1.4583\cdot 10^{-5}}{1+\left(\dfrac{8.4}{G}\right)^{1.7}}-\dfrac{1.7361\cdot 10^{-5}}{1+\left(\dfrac{G}{4.8}\right)^{8.5}}\right),\\ \dot{I}&=\beta\cdot\dfrac{G^{2}}{\alpha^{2}+G^{2}}-\gamma\cdot I,\\ y&=G.\end{array}\right.$ (18)
This model represents the regulation of plasma glucose concentration, $G,$ by means of insulin ($I$), which is secreted by $\beta$ cells. This model has been slightly modified from the one presented by Karin et al. (2016), by removing one parameter ($p$) and transforming one variable ($\beta$), as well as by considering $s_{i}$ as a known constant. While the original model was structurally unidentifiable, the one given by (18), which has three unknown parameters ($c,\alpha,\gamma$), is SLI.
Thus, we have three models ( $\mathcal{M}_{3.2}$, $\mathcal{M}_{3.4}$, and $\mathcal{M}_{3.5}$) that are structurally locally identifiable and observable. Let us now calculate their observability-identifiability matrices, $\mathcal{O}_{I}$, and replace the symbolic variables with their nominal numerical values. The SVD of the resulting arrays yields the singular values plotted in logarithmic scale in Figure 3. None of these plots exhibits a very large gap between singular values, and the largest condition number ($2.2\cdot 10^{6}$) is that of $\mathcal{M}_{3.2}$. However, when we compare their absolute values, we find that the smallest ones are in the order of $10^{-1}$ for the first two examples, while for $\mathcal{M}_{3.5}$ they decrease to $10^{-4}$ and $10^{-5}$. These values suggest that $\mathcal{M}_{3.5}$ has practical identifiability issues.
Figure 3: Singular values of models $\mathcal{M}_{3.2}$, $\mathcal{M}_{3.4}$, and $\mathcal{M}_{3.5}$, in logarithmic scale.
To test this prediction, we explore the practical identifiability of this model by performing a bootstrap procedure, as in the previous cases. Simulating the model with nominal values, we generate 100 pseudo-experimental realizations adding Gaussian noise with zero mean and 0.2 standard deviation, and we solve the 100 optimization problems. Figure 4 shows the fit obtained for one realization; the others were similar. Figure 5 plots the histograms of bootstrap solutions for all the estimated variables: three unknown parameters, $c,\alpha,\gamma,$ and the three initial conditions, $G(t=0),I(t=0),\beta(t=0)$. The values of two of the parameters, $c$ and $\alpha,$ and of the initial condition of the measured variable, $G,$ are well estimated. In contrast, parameter $\gamma$ is clearly practically unidentifiable, and the estimates of the remaining initial conditions are highly uncertain too. Thus, the SVD of $\mathcal{O}_{I}$ correctly reports the existence of practical identifiability issues. The SVD also gives correct results for the previous two case studies, which we found were structurally and practically identifiable.
Figure 4: Model $\mathcal{M}_{3.5}$. Noisy data (circles) of the experiment used for parameter estimation, as well as the noiseless simulation of the output, glucose.
Figure 5: Model $\mathcal{M}_{3.5}$. Bootstrap results: histograms of all estimated variables.